5. Consider two firms selling differentiated varieties of a product, e.g., Coke and Pepsi. Each firm j chooses a price pj for its own variety. Since these varieties are close substitutes, the demand that each firm faces depends not only on its own price, but also the price of its competitor. Specifically, the demand for j’s variety is given by Dj (pj , p−j ) = max 0, 60 + p−j − 2pj

Suppose that both firms can produce any amount of their variety at no cost.

(a) Find firm j’s best response function.

(b) Assume that firms choose prices simultaneously and independently. Show that choosing pj = 18 is not rationalizable. [Hint: perform two rounds of iterated dominance]

(c) Assume that firms choose prices simultaneously and independently. Find the unique Nash equilibirum of the game. [Hint: symmetric games typically have a symmetric equilibrium]

(d) Assume that firm 1 chooses its price first, and firm 2 chooses its price second after seeing firms 1’s price. Find the Stackelberg equilibrium of the game. (e) Compare the equilibrium profits of each firm under part (c), withe their profits on part (d).
Answer #1

V. (a) consider two fioms selling differenticated product Coke and Pepsi The demand for js variety is given by Dj ( Pj , Pj)Assume that firms chooses poices simultaneously and Independently. 11111 99999999999 9 . firm js best response es Snodai Pjnultaneous y Mence, choosing Pi=18 is not a national decision C) Assume that firms choose brices simultaneously and independei P; = 60+ Pj - 60720 - 80 120) . Onique Nash Equilibrium is when both firms choose price equal to 20. (simultaneously and inopy 75= 7 Pj os of love stod los P.j = 7542 Pj = 150 - 21.42 ~ 2105 7 ( Substitute the value of P-j into the best response fufor firm j Di Pj 20:35 Dj - 60+ 21:42 - 2(20•35) = 81.42 - 40.70 Di = 40.72 borba T = (40.72) (20:35) 70 Tj = 828.652 x X-3 l